OpenStudy (solidicecream):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = one half x - 8. Can someone please explain how to do this.

OpenStudy (anonymous):

two parallel line have the same slope

OpenStudy (anonymous):

so tell me wat is the slope of the line?

OpenStudy (anonymous):

@SolidIceCream ,with this information(slope and the point) u can write the equation

OpenStudy (anonymous):

y = x / 2 - 8 This is in the form of y = mx + c where m is the gradient. Gradient = 1 / 2 Since the lines are parallel, the gradients are the same. Next, use the point given and substitute the values into this formula, y = y1 + m(x - x1) where (x1 , y1) is the point. y = 6 + (x + 2) / 2 2y = 12 + x + 2 2y = x + 14

OpenStudy (solidicecream):

Do you mean I have to do the point-slope formula

OpenStudy (anonymous):

yes,as TURITW wrote

OpenStudy (anonymous):

Yes

OpenStudy (solidicecream):

ok I got it thanks for explaining how to do it.

OpenStudy (anonymous):

@SolidIceCream :)